3 times Jos present age is 6 years less than his fathers present age.5 years ago the father was 7 times as old as his son.how old are they

Answer by Boreal(15235)

3J+6=F
(F-5)=7(J-5)
F=7J-35+5=7J-30
3J+6=7J-30
4J=36
Jos=9 years
His father is 33
-
when Jos was 4
his father was 28
-
Answer--9 and 33.